OpenStudy (anonymous):

cos(pi/5)cos(pi/2)-sin(pi/5)sin(pi/2)

OpenStudy (anonymous):

Look at angle sum formula.

OpenStudy (anonymous):

\[ \cos(x+y) = \cos(x)\cos(y) - \sin(x)\sin(y) \]

OpenStudy (anonymous):

cos(x) * sin(y) = 1/2 * (sin(x+y) - sin(x - y)) sin(x) * sin(y) = 1/2 * (cos(x-y) - cos(x+y))

OpenStudy (anonymous):

Or wait, you can also use the fact that \(\cos(\pi/2) - 0\) and \(\sin(\pi/2) = 1\).

OpenStudy (anonymous):

those product identities I posted should help you with this.